# Semantic Scene Completion Based 3D Traversability Estimation for Off-Road Terrains

### v1

@@ Abstract
Off-road environments present significant challenges for autonomous ground vehicles due to the absence of structured roads and the presence of complex obstacles, such as uneven terrain, vegetation, and occlusions. Traditional perception algorithms, designed primarily for structured environments, often fail under these conditions, leading to inaccurate traversability estimations. In this paper, ORDformer, a novel multimodal method that combines LiDAR point clouds with monocular images, is proposed to generate dense traversable occupancy predictions from a forward-facing perspective. By integrating multimodal data, environmental feature extraction is enhanced, which is crucial for accurate occupancy estimation in complex terrains. Furthermore, RELLIS-OCC, a dataset with 3D traversable occupancy annotations, is introduced, incorporating geometric features such as step height, slope, and unevenness. Through a comprehensive analysis of vehicle obstacle-crossing conditions and the incorporation of vehicle body structure constraints, four traversability cost labels are generated: lethal, medium-cost, low-cost, and free. Experimental results demonstrate that ORDformer outperforms existing approaches in 3D traversable area recognition, particularly in off-road environments with irregular geometries and partial occlusions. Specifically, ORDformer achieves over a 20% improvement in scene completion IoU compared to other models. The proposed framework is scalable and adaptable to various vehicle platforms, allowing for adjustments to occupancy grid parameters and the integration of advanced dynamic models for traversability cost estimation.
Note to Practitioners—The motivation of this paper is to help identify areas where vehicles can safely drive in off-road terrains, improving both safety and efficiency in navigation. Current methods typically process semantic and geometric passability features separately, resulting in coarse drivable area recognition. This paper presents ORDformer that combines semantic and geometric features and introduces the RELLIS-OCC dataset specifically for training the model. In real vehicle tests, ORDformer demonstrates effectiveness in challenging off-road scenarios. Future work will integrate ORDformer with 3D path planning algorithms and release datasets with more complex off-road environments to further improve the adaptability and practicality of autonomous navigation systems. This research is applicable to various types of autonomous ground platforms, offering potential for optimizing vehicle autonomy systems and enhancing traversability performance under complex conditions. This can assist engineers in improving the capability of vehicles to navigate challenging environments.
@@
Traversability estimation, semantic scene completion, multimodal perception fusion, off-road navigation.
@@ I Introduction
Autonomous vehicles are increasingly deployed across various environments, including urban, suburban, and off-road settings. While significant advancements have been made in structured environments with well-defined roads and traffic regulations, navigating unstructured off-road terrains remains a substantial challenge. Off-road environments lack clear pathways and present irregular obstacles, as well as diverse terrain features such as uneven ground, vegetation, and occlusions caused by rocky piles or dense foliage, as shown in Fig. 1. These complexities require advanced perception capabilities to ensure safe and efficient navigation.
@@
In urban environments, datasets such as KITTI [1], nuScenes [2], Waymo Open [3], and ApolloScape [4] provide valuable real-world data for obstacle detection and localization. Traditional perception methods [5, 6, 7, 8], designed for 3D object detection tasks, detect and localize obstacles, representing them with bounding boxes. However, these methods struggle to identify irregular and continuously distributed objects in off-road conditions. Recent approaches [9, 10, 11, 12] for 3D occupancy prediction focus on capturing fine-grained spatial structures and semantic information through voxel-based representations. While these methods provide more accurate modeling of obstacles, off-road scenarios also require evaluating the traversability of the environment.
@@
To address these challenges, ORDformer is proposed as a novel multimodal method for off-road traversability estimation. This approach integrates LiDAR point clouds with monocular camera images to generate dense semantic occupancy predictions from a forward-facing perspective. By leveraging the complementary advantages of LiDAR’s geometric precision and the rich semantic information provided by camera sensors, ORDformer enhances the extraction of environmental features, which is essential for accurate occupancy estimation in complex terrains.
@@
Furthermore, RELLIS-OCC, a dataset featuring 3D traversable occupancy annotations, is introduced. This dataset encompasses geometric features such as step height, slope, and unevenness, essential for evaluating the relationship between terrain traversability and vehicle dynamics. Through comprehensive analysis of vehicle obstacle-crossing conditions and incorporation of vehicle body structure constraints, a mapping from voxel semantic categories to traversability cost is established, resulting in four traversability cost labels: lethal, medium-cost, low-cost, and free. This mapping provides a detailed understanding of terrain difficulty, offering precise information for autonomous navigation decision-making.
@@
Building upon this foundation, extensive experiments demonstrate that ORDformer significantly outperforms existing approaches in 3D traversable area recognition, especially in off-road environments where obstacles exhibit irregular geometries and are often partially occluded. Specifically, ORDformer achieves over a 20% improvement in the scene completion IoU metric compared to other models. This performance boost underscores the effectiveness of the multimodal fusion and advanced feature extraction techniques.
@@
Moreover, the proposed framework is highly scalable and adaptable to various vehicle platforms. It offers the flexibility to adjust occupancy grid parameters and can incorporate more sophisticated dynamic models for traversability cost estimation. This adaptability ensures that the method can be tailored to the specific geometric and dynamic characteristics of different autonomous vehicles, enhancing their ability to navigate complex off-road environments safely and efficiently.
@@
TABLE I: Comparison of Different Off-Road Environment Dataset.(Pro.: Proprioceptive, Seg.: Segmentation, Det.: Detection, Exc.: Excitation, SSD: Self-supervised Driving, Ann.: Annotation)
@@
DeepScene [14] 2016 $1024\times 768$checkmark $\checkmark$ - - - - - - $\checkmark$ - - - - - 15,000 366 pixel-wise
YCOR [15] 2018 $1024\times 544$ 1,076 1,076 pixel-wise
RUGD [16] 2019 $1376\times 1110$checkmark - - - - - 37,000 7,546 pixel-wise
Gresenz [17] 2021 $3840\times 2160$checkmark $\checkmark$ - $\checkmark$ - - - - $\checkmark$ - 12,982 - -
RELLIS-3D [18] 2021 $1920\times 1200$checkmark - - 32-line/ 64-line - $\checkmark$ - - - $\checkmark$ $\checkmark$ - $\checkmark$ - - 6,235 6,235 /13,556 pixel-wise /point-wise
ROOAD [19] 2021 $1920\times 1200$checkmark $\checkmark$ - - - - - $\checkmark$ - - 20,000 - -
CaT [20] 2022 $1928\times 1208$checkmark - - - - - 12,300 1,812 traversability
ORFD [21] 2022 $1280\times 720$ 40-line - - - - - $\checkmark$ - - - - - 12,198 12,198 traversability
SORT [22] 2022 $1280\times 720$checkmark - - - - - 20,834 20,834 pixel-wise
TartanDrive [23] 2022 $1024\times 512$checkmark $\checkmark$ - - - $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ - - - $\checkmark$ - $\checkmark$ 5 hours - -
WaterScenes [24] 2023 $1920\times 1080$checkmark $\checkmark$ $\checkmark$ - - $\checkmark$ $\checkmark$ $\checkmark$ - - - 54,120 54,120/ 202,807 /54,120 pixel-wise/ bounding box /point-wise
@@
GOOSE [26] 2024 $2448\times 2048$checkmark 32-line/ 128-line - $\checkmark$ $\checkmark$ - - $\checkmark$ $\checkmark$ - $\checkmark$ - - 15,000 10,000 /10,000 pixel-wise /point-wise
TartanDrive v2.0 [27] 2024 $1024\times 512$checkmark $\checkmark$ - solid/ 32-line - $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ - - - $\checkmark$ $\checkmark$ $\checkmark$ 7 hours - -
@@
    ORDformer, a novel framework utilizing a deformable attention mechanism to integrate LiDAR point clouds and monocular images for 3D traversability prediction, is proposed. To our knowledge, this constitutes the first work to employ semantic scene completion for generating environmental traversable occupancy estimations.
@@
    A pipeline for generating 3D traversable occupancy annotations is introduced, alongside the RELLIS-OCC dataset, a 3D semantic occupancy dataset captured from a forward-facing camera perspective. This dataset addresses the existing gap in 3D traversability annotations for off-road environments.
@@
    ORDformer improves the scene completion IoU metric by over 20% compared to existing models, achieving mean IoU scores of 10.50 and 15.70 for semantic and traversability cost annotations, respectively. In real vehicle experiments, ORDformer performs well in complex off-road scenarios.
The rest of the paper is organized as follows. Section II provides a brief review of the relevant literature. Subsequently, the proposed network and dataset is detailed in Section III. Experimental results and discussion are provided in Section IV. Finally, the conclusion is drawn in Section V.
@@ II-A Off-road Environment Datasets
Perception technology in autonomous driving largely relies on large-scale training datasets; however, existing datasets often do not clearly distinguish between off-road and unstructured road environments. Due to this situation, datasets designed for unstructured roads have been compiled and summarized, as presented in Table I. Public datasets can be categorized based on their data sources into real-vehicle datasets [13, 14, 15, 16, 17, 18, 19, 20, 21, 23, 27, 24, 25, 26] and simulation datasets [22]. Due to differences in data acquisition platform configurations, these datasets typically contain image [14, 15, 16], LiDAR point clouds [13, 25], positioning data [19], or multimodal data [18, 24]. Notably, the TartanDrive [23] dataset, in addition to multimodal data, includes vehicle behavior data (pedal positions, steering) and vehicle proprioceptive data (suspension travel, wheel speed). Some datasets of unstructured roads include annotation information, which can be divided into semantic annotations [16, 18] and traversable area annotations [20, 21]. The RUGD [16] and RELLIS-3D [18] datasets employ the same labeling rules to annotate the semantic categories of objects in the environment. The CaT [20] and ORFD [21] datasets feature traversable area annotations made by experienced drivers, who subjectively judge based on the type of platform.
@@ II-B Traversability Estimation Scheme
Dima et al.[28] defined obstacles as ”areas that a vehicle cannot or should not traverse.” The initiation of the DARPA Challenge in 2004 significantly advanced research on traversable regions in off-road environments[28, 29, 30, 31, 32, 33]. Early studies frequently employed Support Vector Machines (SVMs) to recognize vehicle traversable areas using various perception schemes, including LiDAR and vision fusion[34], pure LiDAR methods[35], and pure vision approaches[36]. With the development of deep learning, these techniques have been widely applied to off-road research, encompassing traversable area image segmentation [37, 38], road detection and classification[39, 40, 41], LiDAR point cloud obstacle recognition[42, 43], multi-sensor fusion[44, 45, 46, 47], and weakly supervised traversable area recognition[48]. Traversable area recognition results are typically represented in various forms, including images[49], 2D grid maps[50], 2.5D grid maps[44], and 3D grid maps[51].
@@ II-C Traversability Quantification
Goodin et al.[52] defines ”traversability” as a quantification of the drivability of vehicles with certain characteristics on terrain. In off-road scenarios, layered cost maps composed of multiple sub-maps addressing different terrain features are widely used for traversable area analysis[53]. These layers typically include ground semantic and geometric information, making the selection of appropriate cost functions crucial for accurate traversability estimation.
Cameras capture rich semantic scene data, while LiDAR provides precise geometric information; thus, fusing these modalities is a common approach[54, 51]. Zhao et al.[54] integrated images and LiDAR point clouds to construct an elevation-semantic 2D grid map, assigning traversability probabilities to various semantic categories to generate a traversability probability grid. Leung et al.[55] developed a traversability cost map by combining image-based semantic segmentation with geometric attributes such as slope, step height, and roughness. Shaban et al.[43] classified point clouds into four traversability levels: free passage, low cost, medium cost, and obstacles, using deep learning networks to produce a Bird’s Eye View (BEV) traversability map.
@@ III Methods
The objective is to predict a dense semantic scene within the camera view in front of the platform by leveraging monocular images and LiDAR point clouds as input, thereby enabling accurate and reliable traversability estimation in off-road environments. Specifically, the image $I_{t}$ and LiDAR point cloud $L_{t}$ are taken as inputs, and the output is a voxel grid $Y_{t}\in\{c_{0},c_{1},c_{2},\cdots,c_{M}\}^{H\times W\times Z}$, where each voxel is classified as either empty $\{c_{0}\}$ or occupied by a specific traversability class within $\{c_{1},c_{2}\cdots,c_{M}\}$. In this context, $M$ denotes the total number of traversability classes, while $H$, $W$, and $Z$ represent the length, width, and height of the voxel grid, respectively. To facilitate this process, RELLIS-OCC, the first occupancy dataset with traversability cost annotations specifically designed for off-road scenarios, is introduced.
III-A Architecture of ORDformer
Off-road environments lack structured roads and traffic regulation information, making the traversability of such terrains closely linked to to the geometric constraints of the platform and necessitating detailed geometric ground analysis. Consequently, there is an increased demand for advanced 3D occupancy prediction capabilities in perception algorithms. To address these challenges, a multimodal semantic scene completion method specifically designed for off-road conditions is proposed. This method leverages LiDAR point clouds and monocular images to generate dense semantic occupancy predictions from a forward-facing perspective. The overall pipeline of ORDformer is illustrated in Fig. 2.
Figure 2: Overall framework of ORDformer.
The model begins by extracting features from both modalities: the LiDAR data, which provides precise geometric information, and images, which contribute rich semantic details. These features are then fused using deformable attention mechanisms to effectively integrate spatial and semantic information. Subsequently, the fused features are processed through deformable self-attention layers to enhance the 3D occupancy representation. Finally, this results in a dense occupancy map highlighting traversable areas.

### v2

# 3DTTNet: Multimodal Fusion-Based 3D Traversable Terrain Modeling for Off-Road Environments

@@ Abstract
Off-road environments remain significant challenges for autonomous ground vehicles, due to the lack of structured roads and the presence of complex obstacles, such as uneven terrain, vegetation, and occlusions. Traditional perception algorithms, primarily designed for structured environments, often fail in unstructured scenarios. In this paper, traversable area recognition is achieved through semantic scene completion. A novel multimodal method, 3DTTNet, is proposed to generate dense traversable terrain estimations by integrating LiDAR point clouds with monocular images from a forward-facing perspective. By integrating multimodal data, environmental feature extraction is strengthened, which is crucial for accurate terrain modeling in complex terrains. Furthermore, RELLIS-OCC, a dataset with 3D traversable annotations, is introduced, incorporating geometric features such as step height, slope, and unevenness. Through a comprehensive analysis of vehicle obstacle-crossing conditions and the incorporation of vehicle body structure constraints, four traversability cost labels are generated: lethal, medium-cost, low-cost, and free. Experimental results demonstrate that 3DTTNet outperforms the comparison approaches in 3D traversable area recognition, particularly in off-road environments with irregular geometries and partial occlusions. Specifically, 3DTTNet achieves a 42% improvement in scene completion IoU compared to other models. The proposed framework is scalable and adaptable to various vehicle platforms, allowing for adjustments to occupancy grid parameters and the integration of advanced dynamic models for traversability cost estimation.
@@
Multimodal fusion, terrain modeling, traversability estimation, semantic scene completion.
@@ I Introduction
Autonomous vehicles are increasingly deployed to various environments, including urban, suburban, and off-road settings. While significant advancements have been made in structured environments with well-defined roads and traffic regulations, navigating unstructured off-road terrains remains a substantial challenge. Off-road environments lack clear pathways and usually come with irregular obstacles, as well as diverse terrain features such as uneven ground, vegetation, and occlusions caused by rocky piles or dense foliage, as shown in Fig. 1. These complexities require advanced perception capabilities to ensure safe and efficient navigation.
@@
In urban environments, datasets such as KITTI [1], nuScenes [2], and Waymo Open [3] provide valuable data for obstacle detection and localization. Traditional perception methods [4],[5],[6] which are designed for 3D object detection tasks, detect and localize obstacles, thereby representing them with bounding boxes. However, these methods struggle to identify irregular and continuously distributed objects in off-road conditions. Recent approaches [7],[8],[9] for 3D occupancy prediction focus on capturing fine-grained spatial structures and semantic information through voxel-based representations. While these methods offer more accurate modeling of obstacles, off-road scenarios also require evaluating the traversability of the environment.
@@
To our knowledge, no work has yet employed semantic scene completion for 3D traversable terrain modeling in off-road scenarios. This gap motivates our novel approach to traversable area recognition. To address these challenges, 3DTTNet, a novel multimodal method for off-road traversability estimation, is proposed. This approach integrates LiDAR point clouds with monocular camera images to generate dense traversable predictions from a forward-facing perspective. By leveraging the complementary advantages of LiDAR’s geometric precision and the rich semantic information provided by camera sensors, 3DTTNet enhances the extraction of environmental features, which is essential for accurate terrain modeling in complex terrains.
@@
Furthermore, RELLIS-OCC is introduced, featuring 3D traversable annotations. This dataset encompasses geometric features such as step height, slope, and unevenness, crucial for evaluating the relationship between terrain traversability and vehicle dynamics. Through comprehensive analysis of vehicle obstacle-crossing conditions and incorporation of vehicle body structure constraints, a mapping from voxel semantic categories to traversability cost is established, resulting in four traversability cost labels: lethal, medium-cost, low-cost, and free. This mapping provides a detailed understanding of terrain difficulty, thus offering precise information for autonomous navigation decision-making.
@@
Building upon this foundation, extensive experiments demonstrate that 3DTTNet significantly outperforms existing approaches in 3D traversable area recognition, especially in off-road environments where obstacles exhibit irregular geometries and are often partially obstructed. Specifically, 3DTTNet achieves a 42% improvement in the scene completion IoU metric compared to other models. This performance boost underscores the effectiveness of the multimodal fusion and advanced feature extraction techniques.
@@
Moreover, the proposed framework is highly scalable and adaptable to various vehicle platforms. It offers a flexibility to adjust grid parameters and can incorporate more sophisticated dynamic models for traversability cost estimation. This adaptability ensures that the method can be tailored to the specific geometric and dynamic characteristics of different autonomous vehicles, further enhancing their ability to navigate complex off-road environments safely and efficiently.
@@
TABLE I: Comparison of Different Off-Road Environment Datasets.(Pro.: Proprioceptive, Seg.: Segmentation, Det.: Detection, Exc.: Excitation, SSD: Self-supervised Driving, Ann.: Annotation)
@@
DeepScene [11] 2016 $1024\times 768$ - $\checkmark$ $\checkmark$ - - - - - - $\checkmark$ - - - - - 15,000 366 pixel-wise
YCOR [12] 2018 $1024\times 544$ - - - - - - - - - - - - - - - 1,076 1,076 pixel-wise
RUGD [13] 2019 $1376\times 1110$ - - - - - - - - - $\checkmark$ - - - - - 37,000 7,546 pixel-wise
Gresenz [14] 2021 $3840\times 2160$ - - - - - $\checkmark$ $\checkmark$ - $\checkmark$ - - - - $\checkmark$ - 12,982 - -
RELLIS-3D [15] 2021 $1920\times 1200$ $\checkmark$ - - 32-line/ 64-line - $\checkmark$ - - - $\checkmark$ $\checkmark$ - $\checkmark$ - - 6,235 6,235 /13,556 pixel-wise /point-wise
ROOAD [16] 2021 $1920\times 1200$ - - - - - $\checkmark$ $\checkmark$ - - - - - $\checkmark$ - - 20,000 - -
CaT [17] 2022 $1928\times 1208$ - - - - - - - - - $\checkmark$ - - - - - 12,300 1,812 traversability
ORFD [18] 2022 $1280\times 720$ - - - 40-line - - - - - $\checkmark$ - - - - - 12,198 12,198 traversability
SORT [19] 2022 $1280\times 720$ - - - - - - - - - $\checkmark$ - - - - - 20,834 20,834 pixel-wise
TartanDrive [20] 2022 $1024\times 512$ $\checkmark$ $\checkmark$ - - - $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ - - - $\checkmark$ - $\checkmark$ 5 hours - -
WaterScenes [21] 2023 $1920\times 1080$ - - - - $\checkmark$ $\checkmark$ $\checkmark$ - - $\checkmark$ $\checkmark$ $\checkmark$ - - - 54,120 54,120/ 202,807 /54,120 pixel-wise/ bounding box /point-wise
@@
GOOSE [23] 2024 $2448\times 2048$ - - $\checkmark$ 32-line/ 128-line - $\checkmark$ $\checkmark$ - - $\checkmark$ $\checkmark$ - $\checkmark$ - - 15,000 10,000 /10,000 pixel-wise /point-wise
TartanDrive v2.0 [24] 2024 $1024\times 512$ $\checkmark$ $\checkmark$ - solid/ 32-line - $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ - - - $\checkmark$ $\checkmark$ $\checkmark$ 7 hours - -
@@
    An innovative framework is proposed for 3D traversable terrain modeling in off-road terrains based on semantic scene completion. This framework enables the completion and inference of occluded regions, a capability that traditional traversability estimation methods cannot achieve.
@@
    3DTTNet, a transformer-based model that integrates multi-sensor features for traversability prediction. By leveraging a deformable attention mechanism, 3DTTNet enhances the interaction between multimodal features.
@@
    A novel pipeline for generating 3D traversable annotations is proposed, integrating semantic accessibility analysis and geometric passability considerations. This method effectively addresses occlusion challenges and facilitates end-to-end model training with quantifiable traversability labels.
The remainder of the paper is organized as follows. Section II provides a brief review of the relevant literature. Subsequently, the proposed network and dataset is detailed in Section III. Experimental results and discussion are demonstrated in Section IV. Finally, a conclusion is drawn in Section V.
@@ II-A Off-road Environment Datasets
Perception technology in autonomous driving largely relies on large-scale training datasets; however, existing datasets often do not clearly distinguish between off-road and unstructured road environments. In view of this situation, datasets designed for unstructured roads have been compiled and summarized, as presented in Table I. Public datasets can be categorized based on their data sources into real-vehicle datasets [10],[11],[12],[13],[14],[15],[16],[17],[18],[20],[24],[21],[22],[23] and simulation datasets [19]. Due to differences in data acquisition platform configurations, these datasets typically contain image [11],[12],[13], LiDAR point clouds [10],[22], positioning data [16], or multimodal data [15],[21]. Notably, the TartanDrive [20] dataset, in addition to multimodal data, is comprised of vehicle behavior data (pedal positions, steering) and vehicle proprioceptive data (suspension travel, wheel speed). Some datasets of unstructured roads include annotation information, which can be divided into semantic annotations [13],[15] and traversable area annotations [17],[18]. The RUGD [13] and RELLIS-3D [15] datasets employ the same labeling rules to annotate the semantic categories of objects in the environment. The CaT [17] and ORFD [18] datasets feature traversable area annotations made by experienced drivers, who subjectively judge based on the type of platform. However, despite these advances, there is a lack of 3D traversable datasets specifically tailored for off-road scenarios.
@@ II-B Traversability Estimation Scheme
Dima et al. [25] defined obstacles as ”areas that a vehicle cannot or should not traverse.” The initiation of the DARPA Challenge in 2004 advanced research on traversable regions in off-road environments [25],[26],[27]. Early studies frequently employed Support Vector Machines (SVMs) to recognize vehicle traversable areas using various perception schemes, including LiDAR and vision fusion [28], pure LiDAR methods [29], and pure vision approaches [30]. With the development of deep learning, these techniques have been widely applied to off-road research, encompassing traversable area image segmentation [31],[32], road detection and classification [33], LiDAR point cloud obstacle recognition [34],[35], multi-sensor fusion [36],[37], and weakly supervised traversable area recognition [38]. Traversable area recognition results are typically represented in various forms: images [39], 2D grid maps [40], 2.5D grid maps [41], and 3D grid maps [42],[43]. Among these, 3D grid maps are capable of more precisely characterizing the shape and occupancy information of obstacles, making them a more ideal representation form.
@@ II-C Traversability Quantification
Goodin et al. [44] defines ”traversability” as a quantification of the drivability of vehicles with certain characteristics on terrain. In off-road scenarios, layered cost maps are composed of multiple sub-maps to address different terrain features, which are widely used for traversable area analysis [45],[46]. These layers typically consist of ground semantic and geometric information, making the selection of appropriate cost functions crucial for accurate traversability estimation.
Cameras capture rich semantic scene data, while LiDAR provides precise geometric information; thus, fusing these modalities is a common approach [47],[42],[48]. Zhao et al. [47] integrated images and LiDAR point clouds to construct an elevation-semantic 2D grid map, assigning traversability probabilities to various semantic categories to generate a traversability probability grid. Leung et al. [49] developed a traversability cost map by combining image-based semantic segmentation with geometric attributes such as slope, step height, and roughness. Shaban et al. [35] classified point clouds into four traversability levels: free passage, low cost, medium cost, and obstacles, and used deep learning networks to produce a Bird’s Eye View (BEV) traversability map. However, these methods exhibit limited capability in 3D environmental representation and are unable to effectively infer and reconstruct occluded regions.
@@ III Methods
The objective is to predict a 3D traversable grid map within the camera view in front of the platform by leveraging monocular images and LiDAR point clouds as input, thereby enabling accurate and reliable traversability estimation in off-road environments. Specifically, the image $I_{t}$ and LiDAR point cloud $L_{t}$ are taken as inputs, and the output is a voxel grid $Y_{t}\in\{c_{0},c_{1},c_{2},\cdots,c_{M}\}^{H\times W\times Z}$, where each voxel is classified as either empty $\{c_{0}\}$ or occupied by a specific traversability class within $\{c_{1},c_{2}\cdots,c_{M}\}$. In this context, $M$ denotes the total number of traversability classes, while $H$, $W$, and $Z$ represent the length, width, and height of the voxel grid, respectively. To facilitate this process, RELLIS-OCC is introduced as the first occupancy dataset with traversability cost annotations specifically designed for off-road scenarios.
III-A Architecture of 3DTTNet
Off-road environments lack structured roads and traffic regulation information, therefore the traversability of such terrains is closely linked to the geometric constraints of the platform and necessitates detailed geometric ground analysis. Consequently, there is an increased demand for advanced 3D traversability prediction capabilities in perception algorithms. To address these challenges, a multimodal semantic scene completion method specifically designed for off-road conditions is proposed. This method takes advantages of LiDAR point clouds and monocular images to generate dense 3D traversable area predictions from a forward-facing perspective. The overall pipeline of 3DTTNet is illustrated in Fig. 2.
Figure 2: Overall framework of 3DTTNet. The process starts by extracting multimodal features from LiDAR point clouds and monocular images, with LiDAR providing geometric details and images contributing semantic information. These features are then fused using deformable attention mechanisms to enhance spatial and semantic interactions. Following fusion, the refined features are processed through deformable attention layers to generate a detailed 3D traversability map that highlights traversable areas.
The model begins by extracting features from both modalities: the LiDAR data which provides precise geometric information, and images which contribute rich semantic details. These features are then fused using deformable attention mechanisms to effectively integrate spatial and semantic information. Subsequently, the fused features are processed through deformable self-attention layers, on purpose to enhance the 3D occupancy representation. Eventually, it produces a dense 3D grid map highlighting traversable areas.